Fermi National Laboratory


Accelerator Update for February 25, 2005 - February 28, 2005

Friday February 25
The day shift began with Operations monitoring store 4004, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

Operations put the Recycler into shot setup to stash antiprotons at 9:14 AM. All shot setups inhibits beam to Switchyard (SY).

The Recycler stashes were completed by 10:09 AM. Pbar resumed stacking.

Operations resumed sending beam to MCenter at 10:41 AM.

NuMI experimenters requested beam at 11:32 AM.

Operations received run conditions for NuMI at 12:12 PM.

At 12:33 PM, Operations reported that the first batch for slipstacking was missing. After further investigation, it was noted that the beam was lost after injection into the Booster.

At 1:02 PM, two Booster RF stations (BRF17 & 18) tripped off. They both reset okay.

At 1:15 PM, eight Booster RF stations tripped off. An expert was notified.

A Booster expert reported at 2:02 PM that he had found a dipole power supply's (HS13) ramp messing up injection.

At 2:23 PM, BRF10 tripped off on a ground fault indication. A minute later BRF7 tripped off. Another minute later eight more RF stations tripped off. An expert reported that all of these trips were due to ongoing balancer work. This work was completed at 3:18 PM.

At 3:32 PM, cryo system technicians at CHL reported that the Tevatron (TeV) F3 sector wet engine had a failing belt. They will replace the belt after the store ends.

At 10:02 PM, Operations reported that the Linac LE Buncher had gone to zero for the second time in the last thirty minutes.

Saturday February 26
The midnight shift began with Operations monitoring store 4004 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

NuMI experts completed their studies at 1:46 AM.

Operations terminated store 4004 at 6:26 AM.

Operators prepared D-Zero (D0) for an access into their collision hall at 6:37 AM.

Cryo system techs completed their F3 wet engine work at 7:30 AM.

D0 came out of access at 9:52 AM.

Operations put the TeV through a dry squeeze at 10:14 AM.

Operations put the TeV into shot setup at 11:07 AM.

At 11:59 AM, Operations reported that a Booster RF station (BRF4) had tripped off three times and now won't reset. Operators contacted an expert and bypassed the station.

Operations established store 4006 at 2:22 PM with an initial luminosity of 88.53E30. Pbar resumed stacking.

NuMI studiers resumed taking beam at 2:45 PM.

Operations resumed sending beam to MCenter at 3:37 PM.

NuMI studiers completed their studies at 8:32 PM.

Operations readied the Recycler to stash at 8:33 PM.

At 8:55 PM, BRF17 tripped off on a ground fault indication. It reset okay.

At 9:23 PM, BRF16 tripped off. It reset okay.

Pbar resumed stacking at 9:31 PM.

MCenter resumed taking beam at 9:37 PM.

Sunday February 27
The midnight shift began with Operations monitoring store 4006 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 4:56 AM, a Booster beam detector tripped off the Booster beam and wouldn't reset. Operators changed the detector and Pbar resumed stacking at 5:30 AM.

NuMI experts began a study period around 8 AM.

At 11:31 AM, Operations reported that the Booster was having serious beam problems at transition.

At 3:40 PM, Operations reported a Pbar Debuncher kicker (D:EKIK) had tripped off. Operators investigated and then called in and expert. The expert had the kicker repaired by 6:22 PM.

Operations terminated store 4006 at 7 PM and then put the TeV into shot setup at 7:30 PM.

At 8:46 PM, the Main Injector (MI) suffered a vacuum burst. It reset okay.

At 8:57 PM, MI lost its beam permit twice due to MECAR trips. They both reset okay.

At 10:12 PM, Operations reported that BRF17 tripped off for the first of five times in forty-five minutes.

Operations established store 4007 at 10:28 PM with an initial luminosity of 93.47E30. Pbar resumed stacking.

MCenter resumed taking beam at 10:58 PM.

A Recycler expert began a study period at 11:06 PM.

Monday February 28
The midnight shift began with Operations monitoring store 4007, with Pbar stacking, with the Recycler conducting studies, and with MiniBooNE and MCenter taking beam.

At 12:28 AM, Operations reported having problems with two Linac quadrupole power supplies (QPS113 & 114) not regulating properly.

At 12:33 AM, the TeV suffered a sector E1 quench. Experts came in, investigated, and reported that the quench was due to a bad controls repeater at sector A1.

Operations reported at 5:57 AM that experts had replaced the sector A1 repeater.

MCenter asked to have their beam turned off at 7:02 AM.

Operations reported at 7:16 AM that they were starting to turn the TeV back on.

The Plan for Monday
The plan for today is to conduct TeV studies and then establish a new store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/28/2005   email Fermilab